OpenStudy (anonymous):

a table for y=x^2-25 is given below. Solve each equation or inequality. a. x^2-25=0 b. x^2-25<0 c. x^2-25>0 table x= -6, -5, -2, 0, 2, 5, 6 y= 11, 0, -21, -25, -21, 0, 11

OpenStudy (goformit100):

I think c x^2-25> But hey there is a big reason behind it

OpenStudy (anonymous):

just put the values of x and y in equation y=x^2-25

OpenStudy (anonymous):

like for y=11 n x=-6 on left term y=11 on right x^2 - 25 = 36-25=11

OpenStudy (anonymous):

i guess you are supposed to divide values of x and y according to it? put values of x and group them in >0 ,<0 and =0

OpenStudy (anonymous):

like for 1st 1 x=-6 -6^2 - 25 = 11 >0 so -6,11 comes in x^2-25>0 group

OpenStudy (anonymous):

are you following me?

OpenStudy (anonymous):

i got it!!!

OpenStudy (anonymous):

good

OpenStudy (anonymous):

wait... that's not right. cause i have a different answer key. ah! this is so confusing